# Recent and Upcoming Developments in Randomized Numerical Linear Algebra for Machine Learning

### v1

@@ 2.1 Matrix Multiplication
If we let $S$ be an $s\times n$ matrix whose $t^{\mathrm{th}}$ row ($t=1,...,s$) has one non-zero, $S_{j_{t}t}=1/\sqrt{sp_{j_{t}}}$, then we can use a “sampling matrix” formalism [67] (a type of “sketching operator”) to express this as:
$AB\approx\frac{1}{s}\sum_{t=1}^{s}\frac{1}{p_{i_{t}}}a_{i_{t}}b_{i_{t}}^{\scriptscriptstyle{\top}}=\left(AS^{\scriptscriptstyle{\top}}\right)\cdot(SB)=C\cdot R,$
where $a_{i_{t}}$ and $b_{i_{t}}^{\scriptscriptstyle{\top}}$ are corresponding columns of $A$ and rows of $B$. Now, if we use probabilities $p_{i}\propto\|a_{i}\|_{2}\|b_{i}\|_{2}$ (row-norm sampling [67]), then we can minimize a very natural Frobenius norm error:
@@
where $\preceq$ denotes the PSD Loewner ordering. (We use $A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA\approx_{\epsilon}A^{\scriptscriptstyle{\top}}A$ as a shorthand for (2) throughout.) The property $A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA\approx_{\epsilon}A^{\scriptscriptstyle{\top}}A$ implies that any $i^{th}$ eigenvalue of the sketch $A^{\scriptscriptstyle{\top}}SSA$ is an $\epsilon$ approximation of the corresponding eigenvalue of the data matrix $A^{\scriptscriptstyle{\top}}A$ (and similarly for the singular values of $SA$). It also extends to the control of various matrix functions of $A^{\scriptscriptstyle{\top}}A$, including the inverse, i.e., $(A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA)^{-1}\approx_{\epsilon}(A^{\scriptscriptstyle{\top}}A)^{-1}$, square root, trace, quadratic form, and more.
@@ 2.2 Least-squares Approximation
    Sketch-and-Solve: Here ([163, 70, 72]), we randomly construct a smaller LS problem, and then solve it using a traditional NLA method. This is the simplest approach to highlight the structure of RandNLA theory; it is most convenient for TCS-style theory; and it leads to low-precision estimates, e.g., $\epsilon=0.1$.
@@
    Iterative Sketching: Here ([174]), we repeatedly sketch/sub-sample the problem randomly, and then iteratively refine the estimate using a convex optimization method. This includes Preconditioned Weighted SGD, Sketch-and-Project, and Newton Sketch; it is the most convenient for ML-style stochastic optimization; and it leads to moderate precision estimates, e.g., $\epsilon=10^{-3}$.
@@
    Sketch-and-Precondition: Here ([156, 9, 134]), we randomly construct an equivalent but well-conditioned problem, and then solve it using a traditional deterministic NLA iterative method. This is the best (usually) for high-quality numerical solutions; it is most convenient for NLA-style implementations; and it leads to high-precision solutions, e.g., $\epsilon=10^{-10}$.
@@ 2.3 Low-rank approximation
LR matrix approximation via the SVD. Let $A$ denote an $m\times n$ data matrix (for simplicity, suppose that $m\geq n$), and let $V$ denote an “embedding” between the “high” and “low” dimensions; then $\hat{A}=AV$ is a low-dimensional representation of the data, and $\tilde{A}=\hat{A}V^{\scriptscriptstyle{\top}}$ is a LR matrix approximation. If $A$ has rank higher than $k$ (up to $n$), then we can consider the truncated SVD,
@@
$PA=\sum_{i=1}^{n}\sigma_{i}Pu_{i}v_{i}^{\scriptscriptstyle{\top}}=\sum_{i=1}^{k}\sigma_{i}u_{i}v_{i}^{\scriptscriptstyle{\top}}=U_{k}\Sigma_{k}V_{k}^{\scriptscriptstyle{\top}}=A_{k}.$
@@
Column subset selection. There is another problem, the Column subset selection problem (CSSP), that is superficially different than LR matrix approximation via the SVD, but that has a similar underlying linear algebraic structure. The problem is to choose a small number of actual columns from an input matrix $A$ [74, 24, 26]. For most objectives of interest, it is intractable to compute the optimal set of columns. One version of the CSSP is the following: instead of projecting onto singular vectors, one can select a subset of data points, $\{a_{i}:i\in S\}$ for $S\subseteq\{1,...,n\}$, and project onto their span. If we let $P_{S}=VV^{\scriptscriptstyle{\top}}$ be a projection onto the span of $\{a_{i}:i\in S\}$, where $V$ is the embedding matrix, then the problem is the following:
@@
Fundamental Structural Result for LR Approximation. For LR matrix approximation, including both via the SVD and via the CSSP, here is the fundamental structural result that is at the heart of RandNLA-based methods [128]. (As with the fundamental structural result for LS, this is a deterministic result that holds for any matrix $S$, and it is central to RandNLA theory by choosing $S$ to be an appropriate sketching operator.) Given $A\in\mathbb{R}^{m\times n}$, let $V_{k}\in\mathbb{R}^{n\times k}$ be the matrix of the top $k$ right singular vectors of $A$. Let $S\in\mathbb{R}^{n\times r}$ ($r\geq k$) be any matrix s.t. $V_{k}^{\scriptscriptstyle{\top}}S$ has full rank. Then, for any unitarily invariant norm $\xi$,
@@
The basic idea is that $\tilde{U}_{q}$ converges to the span$(U_{k})$, and $\tilde{U}_{q}\tilde{U}_{q}^{\scriptscriptstyle{\top}}A$ gives a good $k$-rank approx of $A$. For more detail, see, e.g., Lemma 4.15 from [170]. For any orthonormal matrix $Z\in\mathbb{R}^{n\times k}$, a basic structural result is that
$\|A-ZZ^{\scriptscriptstyle{\top}}A\|_{2}\leq\|(AA^{\scriptscriptstyle{\top}})^{q}A-ZZ^{\scriptscriptstyle{\top}}(AA^{\scriptscriptstyle{\top}})^{q}A\|_{2}^{\frac{1}{2q+1}}.$
From this, by choosing $q=\mathcal{O}(\log(n)/\epsilon)$, $Z=\tilde{U}_{q}$, and a Gaussian $S$, we can show that with high probability
@@ 3 Foundations of “Modern” RandNLA
In this section, we will describe relatively-recent work that provides a foundation for current and upcoming developments in RandNLA. One surprising aspect of Classical RandNLA theory, as outlined in Sec. 2, is that there is actually very little use of randomness. Essentially, all the randomness is “filtered through” the subspace embedding. That is sufficient for many TCS and NLA style objectives (although with Sketch-and-Solve and Sketch-and-Precondition, the two areas use that embedding differently), but it is often “overkill” when RandNLA methods are used in broader (statistical, ML, optimization) pipelines (that typically use some form of Iterative Sketching). This has applications, in particular, in stochastic optimization and traditional statistical resampling methods, where the JL-style guarantees that come with a subspace embedding are sufficient but not necessary. Modern RandNLA theory basically asks for a stronger form of “Algorithmic Gaussianization” than the JL Lemma provides, in the sense that sketches look more Gaussian-like than what is provided with pair-wise JL-like guarantees. To accomplish this, we must make connections with ideas from nonasymptotic RMT, which have proven fruitful across data science and ML.
@@ 3.1 Algorithmic Gaussianization via RMT and LESS Embeddings
Let us compare/contrast this notion of an $(\epsilon,\delta)$-unbiased estimator with with JL / subspace embeddings in Def. 1. For a subspace embedding, we have that
@@
The premise behind model averaging is that if we can produce a bias-corrected estimator $\tilde{x}$, i.e., $\text{Bias}^{2}(\tilde{x})\ll\text{Var}(\tilde{x})$, then we can boost its accuracy by averaging multiple independent copies of this estimator. When the estimator relies on some linear functional of the inverse matrix $(A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA)^{-1}$, as in sketched LS [81] and a number of other RandNLA algorithms [49], then this follows (with some small additional effort) from (6). For example, for $q$ independent sub-Gaussian sketches $S_{i}$ of size $m\geq C(d+q)$, if we average the LS estimates $\tilde{x}_{i}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|S_{i}(Ax-b)\|_{2}$, then the aggregate estimator $\bar{x}=\frac{1}{q}\sum_{i=1}^{q}\tilde{x}_{i}$ satisfies [81]:
$\mathbb{E}\|A\bar{x}-b\|_{2}\leq(1+\epsilon)\|Ax^{*}-b\|_{2},\quad\text{for}\quad\epsilon=O\Big{(}\frac{d}{qm}\Big{)}.$
@@ 3.2 RMT for Sampling via DPPs
This remarkable connection between DPPs and RandNLA has been exploited not only in LR approximation [62, 94, 48], but also in LS regression [58, 60] and iterative optimization [54, 143, 61], including inversion bias corrections similar to those that are possible for Gaussian and LESS sketches [58, 45]; see [55] for a detailed discussion. One way to explain this success is to view DPPs as a natural non-i.i.d. extension of leverage score sampling. Specifically, it can be shown that the marginal probability of the $i^{th}$ row of $A$ appearing in a DPP sampled set $S$, i.e., $\Pr(i\in S)$, corresponds in a precise sense to the ridge-regularized leverage score of that row [42, 46]. These ridge leverage scores have been introduced (independently of DPPs) as the right sub-sampling technique for LR approximation and related tasks in a long series of RandNLA works, e.g., [3, 140, 161].
@@
Nevertheless, there is still an inevitable computational overhead in using non-i.i.d. sampling methods, such as DPPs, in place of popular i.i.d. sampling techniques such as leverage score and row-norm sampling. This trade-off is justified when the quality and small size of the target sample is of paramount importance, e.g., in active learning, Bayesian optimization, clustering, or coreset selection. However, the RMT techniques for analyzing DPPs have also found applications in developing new guarantees for i.i.d. sampling. In particular, it has been shown recently [61, 50] that after preprocessing matrix $A$ with a randomized Hadamard transform, an i.i.d. uniform sample of $O(k\log k)$ rows with high probability contains a DPP sample of size $k$. This has unlocked new approaches in using RandNLA sampling methods for stochastic optimization, as discussed in Sec. 4.
@@ 4.2 Hessian Sketch
These approaches have also been extended to distributed settings via RMT-based model averaging, with applications in ensemble methods, distributed optimization, and federated learning [111, 110, 144, 49, 93]. Further RandNLA-based Newton-type methods include: Subsampled Newton [76, 158, 19, 18]; Subsampled Hessian approximations via Taylor expansion LiSSA [1] and low-rank approximation [78, 56]; as well as approximating the Hessian diagonal/trace via Hutchinson’s method [136] and Stochastic Lanczos Quadrature, particularly for non-convex problems, e.g., PyHessian [175], AdaHessian [176]; and finally Stochastic Quasi-Newton type methods [107, 137].
@@ 4.3 Sketch-and-Project
The Sketch-and-Project framework has recently gained attention as a powerful methodology within the Iterative Sketching paradigm of RandNLA. While this framework has found applications across the stochastic optimization landscape, it originally arose from randomized algorithms for solving systems of linear equations.
@@
3:     Project the current iterate $x_{t}$ onto the solutions of equation $a_{I_{t}}^{\scriptscriptstyle{\top}}x=b_{I_{t}}$.
@@
This simple procedure has been known for a long time [104], but only with the use of randomization are we able to characterize its convergence [165]: if the Kaczmarz method selects index $I_{t}$ randomly, with probability proportional to $\|a_{I_{t}}\|^{2}$, then:
@@ 5.1 Statistical Learning Approaches
As a key motivating example, let us return to linear regression. So far, we have framed this problem so that given a matrix $A$ and vector $b$, we want to find $x^{*}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|Ax-b\|_{2}^{2}$. However, what if part of this data is withheld, for instance due to the cost of procuring the labels?
Formally, suppose that we are given a dataset with $n$ unlabeled data points: $a_{1},\dots,a_{n}\in\mathbb{R}^{d}$. For each point $a_{i}$, we can decide to query the random label $b_{i}\in\mathbb{R}$. Our goal is to query subset $S$ of at most $k\ll n$ labels and output $\hat{x}$ that has small error
$\mathbb{E}\ f(\hat{x}):=\mathbb{E}\ \frac{1}{n}\sum_{i=1}^{n}(a_{i}^{\scriptscriptstyle{\top}}\hat{x}-b_{i})^{2}.$
@@
This task might seem dramatically different than the classical LS task that is tackled by Classical RandNLA. Yet, recent works have shown that RandNLA subsampling techniques, such as leverage score sampling and DPPs, play a crucial role in ensuring good generalization error in the setting. In particular, it is known (e.g., [124, 60]) that for any fixed $A$ and random $b$, if we query $O(d\log d+d/\epsilon)$ entries of $b$ sampled according to $A$’s leverage scores, and compute $\hat{x}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|A_{S}x-b_{S}\|_{2}^{2}$, then:
@@
RandNLA subsampling methods have also been extended to “non-linear” learning tasks, via kernel-based learning methods. These approaches work by constructing a mapping $\phi:\mathcal{X}\rightarrow\mathcal{F}$ from the data space $\mathcal{X}$ to an expanded feature space $\mathcal{F}$. While the expanded data vectors are never explicitly computed, we can learn about the geometry of the dataset by looking at the pairwise inner products between $i^{th}$ and $j^{th}$ data point, which are obtained through the so-called kernel funtion $k(x_{i},x_{j})=\phi(x_{i})^{\scriptscriptstyle{\top}}\phi(x_{j})$. In the resulting learning problem, we receive a dataset $(x_{i},y_{i})_{i=1}^{n}\sim\mathcal{D}$, and construct an estimator $\tilde{f}$ for $f^{*}$ by minimizing a regularized objective over the kernel space $\mathcal{F}$:
$\min_{f\in\mathcal{F}}\ \frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i}),y_{i})+\frac{\lambda}{2}\|f\|_{\mathcal{F}}^{2}.$
This is equivalent to solving the following regularized linear system:
@@
Naturally, we could now use RandNLA techniques, such as those discussed in Sec. 4, to solve (10) directly. However, this being a statistical learning problem significantly affects our algorithmic choices. To start, note that we are not given $K$ as input, but rather we must compute each of its entries via the kernel function $k(\cdot,\cdot)$. In fact, even just forming $K$ at the cost of $O(n^{2})$ kernel evaluations is a computational bottleneck we may wish to avoid. On the other hand, due to the presence of generalization error, we do not need to solve the problem to high precision.
We can address both of these considerations by exploiting the fact that the kernel matrix $K$ has a PSD structure. This allows us to approximate it efficiently without even forming the entire matrix $K$ by relying on the so-called Nyström approximation:
$\tilde{K}$ $=KS^{\scriptscriptstyle{\top}}(SKS^{\scriptscriptstyle{\top}})^{-1}SK,$
where $S$ is a $k\times n$ subsampling matrix such that $KS^{\scriptscriptstyle{\top}}$ selects $k$ columns from $K$ and $SK$ selects $k$ rows from $K$. The Nyström method requires only $nk$ kernel evaluations, where $k$ is a rank parameter that controls the accuracy. A long line of works at the intersection of ML and RandNLA have shown that this method can be used to produce kernel ridge regression estimators with strong generalization guarantees. While early efforts proposed using uniform subsampling to select the rows/columns of $K$ [169, 11, 162], more recent works have shown that RandNLA-style subsampling using the so-called ridge leverage scores yields more robust estimators [3, 140, 161]. This has been further improved with DPP sampling [46, 29, 48].
Further related works on the Nyström method include applications to Gaussian Process regression [29, 30], Nyström-based sketching [87, 15] and preconditioning [8, 79, 78, 56], sublinear time LR approximation of other structured matrices [142, 14], and sublinear time DPP sampling algorithms [46, 31, 7].

### v2

@@ 2.1 Matrix Multiplication
If we let $S$ be an $s\times n$ matrix whose $t^{\mathrm{th}}$ row ($t=1,...,s$) has one non-zero, $S_{t,j_{t}}=1/\sqrt{sp_{j_{t}}}$, then we can use a “sampling matrix” formalism [66] (a type of “sketching operator”) to express this as:
$AB\approx\frac{1}{s}\sum_{t=1}^{s}\frac{1}{p_{j_{t}}}a_{j_{t}}b_{j_{t}}^{\scriptscriptstyle{\top}}=\left(AS^{\scriptscriptstyle{\top}}\right)\cdot(SB)=C\cdot R,$
where $a_{j_{t}}$ and $b_{j_{t}}^{\scriptscriptstyle{\top}}$ are corresponding columns of $A$ and rows of $B$. Now, if we use probabilities $p_{i}\propto\|a_{i}\|_{2}\|b_{i}\|_{2}$ (row-norm sampling [66]), then we can minimize a very natural Frobenius norm error:
@@
where $\preceq$ denotes the PSD Loewner ordering. (We use $A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA\approx_{\epsilon}A^{\scriptscriptstyle{\top}}A$ as a shorthand for (2) throughout.) The property $A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA\approx_{\epsilon}A^{\scriptscriptstyle{\top}}A$ implies that any $i^{\mathrm{th}}$ eigenvalue of the sketch $A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA$ is an $\epsilon$ approximation of the corresponding eigenvalue of the data matrix $A^{\scriptscriptstyle{\top}}A$ (and similarly for the singular values of $SA$). It also extends to the control of various matrix functions of $A^{\scriptscriptstyle{\top}}A$, including the inverse, i.e., $(A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA)^{-1}\approx_{\epsilon}(A^{\scriptscriptstyle{\top}}A)^{-1}$, square root, trace, quadratic form, and more.
@@ 2.2 Least-squares Approximation
    Sketch-and-Solve: Here [164, 69, 71], we randomly construct a smaller LS problem, and then solve it using a traditional NLA method. This is the simplest approach to highlight the structure of RandNLA theory; it is most convenient for TCS-style theory; and it leads to low-precision estimates, e.g., $\epsilon=0.1$.
@@
    Iterative Sketching: Here [91, 151, 175], we repeatedly sketch/sub-sample the problem randomly, and then iteratively refine the estimate using a convex optimization method. This includes Preconditioned Weighted SGD, Sketch-and-Project, and Newton Sketch; it is the most convenient for ML-style stochastic optimization; and it leads to moderate precision estimates, e.g., $\epsilon=10^{-3}$.
@@
    Sketch-and-Precondition: Here [157, 9, 134], we randomly construct an equivalent but well-conditioned problem, and then solve it using a traditional deterministic NLA iterative method. This is the best (usually) for high-quality numerical solutions; it is most convenient for NLA-style implementations; and it leads to high-precision solutions, e.g., $\epsilon=10^{-10}$.
@@ 2.3 Low-rank approximation
LR matrix approximation via the SVD. Let $A$ denote an $m\times n$ data matrix (for simplicity, suppose that $m\geq n$), and let $Q\in\mathbb{R}^{m\times k}$ denote an “embedding” between the “high” and “low” dimensions; then $\hat{A}=Q^{\scriptscriptstyle{\top}}A$ is a low-dimensional representation of the data, and $\tilde{A}=QQ^{\scriptscriptstyle{\top}}A$ is a LR matrix approximation. If $A$ has rank higher than $k$, then the truncated SVD give a LR approximation,
@@
$PA=U_{k}U_{k}^{\scriptscriptstyle{\top}}A=\sum_{i=1}^{n}\sigma_{i}U_{k}U_{k}^{\scriptscriptstyle{\top}}u_{i}v_{i}^{\scriptscriptstyle{\top}}=\sum_{i=1}^{k}\sigma_{i}u_{i}v_{i}^{\scriptscriptstyle{\top}}=U_{k}\Sigma_{k}V_{k}^{\scriptscriptstyle{\top}}=A_{k}.$
@@
Column subset selection. There is another problem, the Column subset selection problem (CSSP), that is superficially different than LR matrix approximation via the SVD, but that has a similar underlying linear algebraic structure. The problem is to choose a small number of actual columns from an input matrix $A$ [73, 23, 25]. For most objectives of interest, it is intractable to compute the optimal set of columns. One version of the CSSP is the following: instead of projecting onto singular vectors, one can select a subset of data points, $\{a_{i}:i\in S\}$ for $S\subseteq\{1,...,n\}$, and project onto their span (with slight abuse of notation, here $S$ is a subset, while later it is a sketching operator). If we let $P_{S}=QQ^{\scriptscriptstyle{\top}}$ be a projection onto the span of $\{a_{i}:i\in S\}$, where $Q$ is the embedding matrix, then the problem is the following:
@@
Fundamental Structural Result for LR Approximation. For LR matrix approximation, including both via the SVD and via the CSSP, here is the fundamental structural result that is at the heart of RandNLA-based methods [128]. (As with the fundamental structural result for LS, this is a deterministic result that holds for any matrix $S$, and it is central to RandNLA theory by choosing $S$ to be an appropriate sketching operator.) Given $A\in\mathbb{R}^{m\times n}$, let $V_{k}\in\mathbb{R}^{n\times k}$ be the matrix of the top $k$ right singular vectors of $A$. Let $S\in\mathbb{R}^{n\times s}$ ($s\geq k$) be any matrix s.t. $V_{k}^{\scriptscriptstyle{\top}}S$ has full rank. Then, for any unitarily invariant norm $\xi$,
@@
The basic idea is that $\tilde{U}_{q}$ converges to the span$(U_{k})$, and $\tilde{U}_{q}\tilde{U}_{q}^{\scriptscriptstyle{\top}}A$ gives a good $k$-rank approx of $A$. For more detail, see, e.g., Lemma 4.15 from [171]. For any orthonormal matrix $Q\in\mathbb{R}^{m\times k}$, a basic structural result is that
$\|A-QQ^{\scriptscriptstyle{\top}}A\|_{2}\leq\|(AA^{\scriptscriptstyle{\top}})^{q}A-QQ^{\scriptscriptstyle{\top}}(AA^{\scriptscriptstyle{\top}})^{q}A\|_{2}^{\frac{1}{2q+1}}.$
From this, by choosing $q=\mathcal{O}(\log(n)/\epsilon)$, $Q=\tilde{U}_{q}$, and a Gaussian $S$, we can show that with high probability
@@ 3 Foundations of “Modern” RandNLA
In this section, we will describe relatively-recent work that provides a foundation for current and upcoming developments in RandNLA. One surprising aspect of Classical RandNLA theory, as outlined in Sec. 2, is that there is actually very little use of randomness. Essentially, all the randomness is “filtered through” the subspace embedding. That is sufficient for many TCS and NLA style objectives (although with Sketch-and-Solve and Sketch-and-Precondition, the two areas use that embedding differently), but it is often “overkill” when RandNLA methods are used in broader (statistical, ML, optimization) pipelines (that typically use some form of Iterative Sketching). This has applications, in particular, in stochastic optimization and traditional statistical resampling methods, where the JL-style guarantees that come with a subspace embedding are sufficient but not necessary. Modern RandNLA theory basically asks for a stronger form of “Algorithmic Gaussianization” than the JL Lemma provides, in the sense that sketches look more Gaussian-like than what is provided with pair-wise JL-like guarantees. To accomplish this, we must make connections with ideas from non-asymptotic RMT, which have proven fruitful across data science and ML.
@@ 3.1 Algorithmic Gaussianization via RMT and LESS Embeddings
Let us compare/contrast this notion of an $(\epsilon,\delta)$-unbiased estimator with JL / subspace embeddings in Def. 1. For a subspace embedding, we have that
@@
The premise behind model averaging is that if we can produce a bias-corrected estimator $\tilde{x}$, i.e., $\text{Bias}^{2}(\tilde{x})\ll\text{Var}(\tilde{x})$, then we can boost its accuracy by averaging multiple independent copies of this estimator. When the estimator relies on some linear functional of the inverse matrix $(A^{\scriptscriptstyle{\top}}S^{\scriptscriptstyle{\top}}SA)^{-1}$, as in sketched LS [80] and a number of other RandNLA algorithms [48], then this follows (with some small additional effort) from (6). For example, for $q$ independent sub-Gaussian sketches $S_{i}$ of size $s\geq O(d+q)$, if we average the LS estimates $\tilde{x}_{i}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|S_{i}(Ax-b)\|_{2}$, then the aggregate estimator $\bar{x}=\frac{1}{q}\sum_{i=1}^{q}\tilde{x}_{i}$ satisfies [80]:
$\mathbb{E}\|A\bar{x}-b\|_{2}\leq(1+\epsilon)\|Ax^{*}-b\|_{2},\quad\text{for}\quad\epsilon=O\Big{(}\frac{d}{qs}\Big{)}.$
@@ 3.2 RMT for Sampling via DPPs
This remarkable connection between DPPs and RandNLA has been exploited not only in LR approximation [61, 93, 47], but also in LS regression [57, 59] and iterative optimization [53, 143, 60], including inversion bias corrections similar to those that are possible for Gaussian and LESS sketches [57, 44]; see [54] for a detailed discussion. One way to explain this success is to view DPPs as a natural non-i.i.d. extension of leverage score sampling. Specifically, it can be shown that the marginal probability of the $i^{\mathrm{th}}$ row of $A$ appearing in a DPP sampled set $S$, i.e., $\Pr(i\in S)$, corresponds in a precise sense to the ridge-regularized leverage score of that row [41, 45]. These ridge leverage scores have been introduced (independently of DPPs) as the right sub-sampling technique for LR approximation and related tasks in a long series of RandNLA works, e.g., [3, 140, 162].
@@
Nevertheless, there is still an inevitable computational overhead in using non-i.i.d. sampling methods, such as DPPs, in place of popular i.i.d. sampling techniques such as leverage score and row-norm sampling. This trade-off is justified when the quality and small size of the target sample is of paramount importance, e.g., in active learning, Bayesian optimization, clustering, or coreset selection. However, the RMT techniques for analyzing DPPs have also found applications in developing new guarantees for i.i.d. sampling. In particular, it has been shown recently [60, 49] that after preprocessing matrix $A$ with a randomized Hadamard transform, an i.i.d. uniform sample of $O(k\log k)$ rows with high probability contains a DPP sample of size $k$. This has unlocked new approaches in using RandNLA sampling methods for stochastic optimization methods such as Sketch-and-Project (see Sec. 4).
@@ 4.2 Hessian Sketch
These approaches have also been extended to distributed settings via RMT-based model averaging, with applications in ensemble methods, distributed optimization, and federated learning [110, 109, 144, 48, 92]. Further RandNLA-based Newton-type methods include: Subsampled Newton [75, 159, 18, 17]; Hessian approximations via randomized Taylor expansion [1] and low-rank approximation [77, 55]; Hessian diagonal/trace estimates via Hutchinson’s method [136] and Stochastic Lanczos Quadrature, particularly for non-convex problems, e.g., PyHessian [176], AdaHessian [177]; and finally Stochastic Quasi-Newton type methods [106, 137].
@@ 4.3 Sketch-and-Project
The Sketch-and-Project framework has gained attention as a powerful methodology within the Iterative Sketching paradigm of RandNLA. While this framework has found applications across the stochastic optimization landscape, it originally arose from randomized algorithms for solving systems of linear equations.
@@
3:     $x_{t+1}\leftarrow$ Project $x_{t}$ onto the solutions of equation $a_{I_{t}}^{\scriptscriptstyle{\top}}x=b_{I_{t}}$
@@
This simple procedure has been known for a long time [103], but only with the use of randomization are we able to characterize its convergence [166]: if the Kaczmarz method selects index $I_{t}$ randomly, with probability proportional to $\|a_{I_{t}}\|_{2}^{2}$, then:
@@ 5.1 Statistical Learning Approaches
As a key motivating example, let us return to LS regression. So far, we have framed this problem so that given a matrix $A$ and vector $b$, we want to find $x^{*}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|Ax-b\|_{2}^{2}$. However, what if part of this data is withheld, for instance due to the cost of procuring the labels?
Formally, suppose that we are given a data matrix $A\in\mathbb{R}^{n\times d}$ with $n$ unlabeled data points: $a_{1},\dots,a_{n}\in\mathbb{R}^{d}$ (transposed rows of $A$). For each point $a_{i}$, we can decide to query the random label $b_{i}\in\mathbb{R}$. Our goal is to query subset $S\subseteq\{1,...,n\}$ of at most $s\ll n$ labels and output $\hat{x}$ that has small error
$\mathbb{E}\,f(\hat{x}):=\mathbb{E}\,\frac{1}{n}\sum_{i=1}^{n}(a_{i}^{\scriptscriptstyle{\top}}\hat{x}-b_{i})^{2}=\mathbb{E}\,\frac{1}{n}\|A\hat{x}-b\|_{2}^{2}.$
@@
This task might seem dramatically different than the classical LS task that is tackled by Classical RandNLA. Yet, recent works have shown that RandNLA subsampling techniques, such as leverage score sampling and DPPs, play a crucial role in ensuring good generalization error in this setting. In particular, it is known (e.g., [124, 59]) that for any fixed $A$ and random $b$, if we query a subset $S$ of $s=O(d\log d+d/\epsilon)$ entries of $b$ sampled according to $A$’s leverage scores, and compute $\hat{x}=\operatorname*{\mathop{\mathrm{argmin}}}_{x}\|A_{S}x-b_{S}\|_{2}^{2}$, then:
@@
RandNLA subsampling methods have also been extended to “non-linear” learning tasks, via kernel-based learning methods. These approaches work by constructing a mapping $\phi:\mathcal{X}\rightarrow\mathcal{F}$ from the data space $\mathcal{X}$ to an expanded feature space $\mathcal{F}$. While the expanded data vectors are never explicitly computed, we can learn about the geometry of the dataset by looking at the pairwise inner products between $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ data point, which are obtained through the so-called kernel function $k(x_{i},x_{j})=\phi(x_{i})^{\scriptscriptstyle{\top}}\phi(x_{j})$. In the resulting learning problem, we receive a dataset $(x_{i},y_{i})_{i=1}^{n}\sim\mathcal{D}$, and construct an estimator $\tilde{f}$ for $f^{*}$ by minimizing a regularized objective over the kernel space $\mathcal{F}$:
$\min_{f\in\mathcal{F}}\ \frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i}),y_{i})+\frac{\lambda}{2}\|f\|_{\mathcal{F}}^{2},$
where $\|f\|_{\mathcal{F}}$ is the norm induced by $\mathcal{F}$ and $\ell$ is a loss function. Using square loss and the representer theorem, this can be cast as kernel ridge regression (KRR), which is equivalent to solving the following regularized linear system:
@@
Naturally, we could now use RandNLA techniques, such as those discussed in Sec. 4, to solve (10) directly. However, this being a statistical learning problem significantly affects our algorithmic choices. To start, note that we are not given $K$ as input, but rather we must compute each of its entries via the kernel function $k(\cdot,\cdot)$. In fact, even just forming $K$ at the cost of $n^{2}$ kernel evaluations is a computational bottleneck we may wish to avoid. On the other hand, due to the presence of generalization error, we do not need to solve the problem to high precision.
We can address both of these considerations by exploiting the fact that the kernel matrix $K$ has a PSD structure. This allows us to approximate it efficiently without even forming the entire matrix $K$ by relying on a Nyström approximation:
$\tilde{K}$ $=KS(S^{\scriptscriptstyle{\top}}KS)^{-1}S^{\scriptscriptstyle{\top}}K,$
where $S$ is an $n\times s$ subsampling matrix such that $KS$ selects $s$ columns from $K$ and $S^{\scriptscriptstyle{\top}}KS$ selects an $s\times s$ principal submatrix of $K$. Thus, the Nyström method requires only $ns$ kernel evaluations, i.e., sublinear in the size of $K$. This approach is closely linked to the LR approximations described in Sec. 2.3: using the PSD property of the kernel matrix $K$, we can express it as $K=A^{\scriptscriptstyle{\top}}A$ for some matrix $A$; then, the Nyström approximation is given by $\tilde{K}=A^{\scriptscriptstyle{\top}}P_{AS}A$, and its spectral/Frobenius error is $\|K-\tilde{K}\|_{\xi}=\|A-P_{AS}A\|_{\xi}^{2}$, which corresponds to the LR projection error (3) for $A$.
A long line of works at the intersection of ML and RandNLA have shown that this method can be used to produce kernel ridge regression estimators with strong generalization guarantees faster than it takes to form the kernel matrix $K$. While early efforts proposed using uniform subsampling to select the rows/columns of $K$ [170, 11, 163], more recent works have shown that RandNLA-style subsampling using ridge leverage scores yields more robust estimators [3, 140, 162]. This has been further improved with DPP sampling [45, 28, 47].
Additional related works on the Nyström method include applications to Gaussian Process regression [28, 29], Nyström-based sketching [86, 14] and preconditioning [8, 78, 77, 55], sublinear time LR approximation of other structured matrices [142, 13], and sublinear time DPP sampling algorithms [45, 30, 7].